OpenStudy (anonymous):

There are three types of razor blades that come in packages of 6, 12, and 24, costing $2, $3, and $4 respectively. The store sold 12 packages containing 162 razor blades altogether for a total cost of $35. How many of each type of package were sold?

OpenStudy (anonymous):

looks to be 4,5, and 3 if I mess with it in my head. But hoping someone could show me in formula

OpenStudy (anonymous):

2x + 3y + 4z = 35 (equation for price) and I think you need to make other equations regarding to number of packages

OpenStudy (anonymous):

for packages I would assume 6x+12y+24z=162 then solve system of solutions

OpenStudy (anonymous):

ding ding ding!!!

OpenStudy (anonymous):

wish I could have thought like this when taking my final an hour ago :(

OpenStudy (anonymous):

so that gets you y=19-7.67z and substitue in?

OpenStudy (anonymous):

hmmm now Im stuck

OpenStudy (anonymous):

Got it... 1,7,3